OpenStudy (anonymous):

4(x-2)/x-3+3/x=-3/x(x-3)

Nnesha (nnesha):

 isn't that you question

Nnesha (nnesha):

idistribute x-2 by 4 frst of all

OpenStudy (campbell_st):

multiply every term by x(x -3) and cancelling common factors you get \[4(x -2)\times x + 3(x -2) = -3\] now distribute and collect like terms \[4x^2 - 5x - 3 = 0\] now solve

OpenStudy (campbell_st):

oops slight error \[4x(x - 2) + 3(x -3) = -3\] so its \[4x^2 - 5x -6 = 0\] now solve

OpenStudy (campbell_st):

the quadratic factors... so you can find the solutions by factoring... or using the general quadratic formula...